Problem 1-36 (LO 1-3) Campbell, a single taxpayer, earns $190,400 in taxable income and $9,500 in interest from an investment in State of New York bonds. (Use the U.S. tax rate schedule.) Required: How much federal tax will she owe? What is her average tax rate? What is her effective tax rate? What is her current marginal tax rate?
Answer #1

Part A

Federal tax = 32748.50+(32%*(190400-160725)) = $42244.50 = $42245

Part B

Average tax = total tax / total taxable income = 42245/190400 = 22.19%

Part C

Effective tax rate = total tax / total income = 42245/(190400+9500) = 21.13%

Part D

Current marginal tax rate = 32%